OpenStudy (anonymous):

You throw a basketball whose path can be modeled by y=-16x^2+15x+6, where (in seconds) and y represents height of the basketball (in feet). A. What is the maximum height that the basketball reaches? B. And how many seconds will the basketball hit the ground if no one catches it?

OpenStudy (anonymous):

I will give any person that can help a medal

OpenStudy (anonymous):

PLEASE HEEEEELP

OpenStudy (anonymous):

we can do this

OpenStudy (anonymous):

\[ y=-16x^2+15x+6\] the maximum height is at the vertex the first coodinate of the vertex is \(-\frac{b}{2a}\) which in this case is\[-\frac{15}{2\times -16}=\frac{15}{32}\]

OpenStudy (anonymous):

Thank you so much for the reply, I have been waiting a while to get the answer. And for that reason I have given you a medal. :)

OpenStudy (anonymous):

Thanks again!!!!!